OpenStudy (anonymous):

a football is thrown in the air.. the height, h(t), of the ball, in metres, after t seconds is modelled by this equation...

OpenStudy (anonymous):

\[h(t) = -4.9(t-1.25)^{2} + 9\]

OpenStudy (anonymous):

how high off the ground was the ball when it was thrown?

OpenStudy (anonymous):

maximum height is 9 after 1.25 seconds

OpenStudy (anonymous):

how did you find the maximum height?

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

easy. the first part is a square. anything squared is positive. then when you multiply by -4.9 is will be negative, so the biggest the first part can be is 0. therefore the biggest the whole thing can be is 9

OpenStudy (anonymous):

so, how high off the ground was the ball when it was thrown ?

OpenStudy (anonymous):

plug in time

OpenStudy (anonymous):

set derivative to 0 solve for t use t to find max height

OpenStudy (anonymous):

answer to that is to expand and see what the constant is

OpenStudy (anonymous):

you get 1.34375

OpenStudy (anonymous):

like heck you do. you do not take the derivative and you do not set it to zero. you just see that the maximum is 9 because it is displayed for you

OpenStudy (anonymous):

okay, how high was the ball at 2.5 sec... (i got 1.3 metres) and was the ball on its way up or down, how do you know?

OpenStudy (anonymous):

maximum height is 9 yes?

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

and we know that it is 9 when x = 1.25 yes?

OpenStudy (anonymous):

i believe soo..

OpenStudy (anonymous):

which means that it is the highest after 1.25 seconds so after that it is heading down

OpenStudy (anonymous):

goes up to 9 at 1.25 seconds, comes back down after

OpenStudy (anonymous):

okay, makes sense now.. last question! when does the ball hit the ground?

OpenStudy (anonymous):

set = 0 and solve for t, because when it hits the ground the height is 0

OpenStudy (anonymous):

okay, thanks :]

OpenStudy (anonymous):

\[-4.9(t-1.25)^2+9=0\] \[(t-1.25)^2=\frac{9}{4.9}\] \[x-1.25=\sqrt{\frac{9}{4.9}}\] \[x=1.25+\sqrt{\frac{9}{4.9}}\]

OpenStudy (anonymous):

whatever that is